# Counterfactual Sensitivity in Quantitative Spatial Models

### v2

@@ 4.3 Results
Having obtained a posterior for the true trade flows given the noisy trade flows, we can now find credible sets for the counterfactual change variables of interest. In Figure 1, I reproduce Figure 3 of Adao, Costinot, and Donaldson (2017), which plots the change in China’s welfare as a result of the China shock for each year in the period 1996-2011, and include 95% equal-tailed confidence intervals or credible sets. Specifically, I plot three different intervals.
The first is a frequentist confidence interval that only considers estimation error and hence assumes the data are perfectly measured. It is constructed using code provided by the authors, which uses a bootstrap procedure that samples from the asymptotic distribution of the GMM estimator for the trade elasticity $\varepsilon$. The resulting confidence intervals are small for the period before the year 2000, and then slowly become wider. For the period 2003-2011, with the exception of 2009, it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
The second region considers only measurement error and not sampling error in $\varepsilon$, so the goal is to find a 95% credible set for $\hat{W}_{\mathrm{China},t}^{E}$ for $t=1996,...,2011$. Towards this, I feed forward draws from the posterior in Equation (17) into the function $g_{\hat{W}_{\mathrm{China},t}}$ in Equation (15). The resulting credible set is slightly wider than the confidence interval based solely on estimation error, especially in the first few years. Again, for the period 2003-2011 it is unclear whether joining the WTO had a positive or negative effect on China’s welfare.
Finally, the third region combines estimation error and measurement error. As outlined in Remark 1, the focus then shifts to finding 97.5% credible sets for the upper and lower endpoints of a 97.5% frequentist confidence interval that considers only estimation error. The resulting lower bound seems to be a reasonable composition of the individual lower bounds considering only estimation error or only measurement error. On the other hand, the posterior distributions of the upper bounds have heavy right tails and the upper bounds become large.
@@
We can learn more from the posterior distributions than just credible sets. It might for example be of interest what the posterior probability is that the welfare in China increased as a result of the China shock, which is calculated in Appendix D.3. There, I also perform some additional analyses to check the robustness of these results.
@@ 5.1 Model and Counterfactual Question of Interest
The empirical application in Allen and Arkolakis (2022) aims to answer the question what the returns on investment are of all the highway segments of the US Interstate Highway network. The authors do so by introducing an economic geography model and calculating what happens to welfare after a small ($1\%$) improvement to all highway links. Combining these counterfactual welfare changes with how many lane-miles must be added in order to achieve the $1\%$ improvement, they find the highway segments with the greatest return on investment.
This exercise requires only data on incomes and traffic flows of the $n$ locations and knowledge of four structural model parameters. Three of these parameters are taken from the literature and assumed to have no uncertainty around them. The fourth, which is the strength of traffic congestion $\delta$, is estimated using the noisily measured traffic flow data. The details of the model can be found in Appendix E.1, but the key relation is the one that maps the average annual daily traffic (AADT) flows $\left\{F_{ij}\right\}$ to estimated welfare $\hat{W}^{E}$, which is
$\hat{W}^{E}=g_{\hat{W}}\left(\left\{F_{ij}\right\};\tilde{\delta}\left(\left\{F_{ij}\right\}\right)\right)$
for known functions $\tilde{\delta}:\mathbb{R}_{+}^{n\left(n-1\right)}\rightarrow\mathbb{R}$ and $g_{\hat{W}}:\mathbb{R}_{+}^{n\left(n-1\right)}\rightarrow\mathbb{R}$.
Similarly as for the application in Section 4, I distinguish between the scenarios where the strength of traffic congestion $\delta$ is estimated with and without estimation error. In Allen and Arkolakis (2022), estimation error is not considered so to obtain the frequentist confidence interval that accounts for estimation error I use the delta method as outlined in Section 2.2.
@@ 5.2 Measurement Error Model and Prior
Towards uncertainty quantification, I will add log-normal measurement error to non-zero traffic flows. Musunuru and Porter (2019) estimates that the measurement error variance of the logarithm of the average annual daily traffic (AADT) flows, which is exactly the data that Allen and Arkolakis (2022) uses, is between 0.05 and 0.20. To obtain a lower bound on uncertainty, I will use a uniform measurement error variance of $0.05$. I follow Musunuru and Porter (2019) by adding independent and identically distributed measurement error shocks to log traffic flows. Hence, this fits the setting from Section 3.4 and the prior and measurement error model are as in Equation (12).
Using $\hat{\varsigma}^{2}=0.05$, I estimate a prior variance of $\hat{s}^{2}=0.101$. This results in the following estimated posterior for the true traffic flow between country $i$ and $j$, $\log F_{ij}$, given its noisy version $\log\tilde{F}_{ij}$, for $i,j=1,...,n$,
$\mathcal{N}\left(0.669\cdot\log\tilde{F}_{ij}+0.331\cdot\left\{\hat{\beta}\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\},0.033\right).$
@@ 5.3 Results
The counterfactual question of interest is which links have the highest return on investment, and the authors of Allen and Arkolakis (2022) report the top ten links. For exposition, I will focus my analysis on the three best performing links. Similarly to the application in Section 4, I consider scenarios with only estimation error, only measurement error, and both estimation and measurement error. Table 1 shows the $95\%$ equal-tailed confidence intervals and credible sets for the top three links. Accounting only for estimation error results in bounds that are around $50\%$ away from the point estimate, and the bounds resulting from accounting only for measurement error are generally less wide.
@@
Only est. error [5.31, 15.55] [4.69, 14.38] [3.62, 11.00]
Only meas. error, EB [8.52, 14.05] [7.42, 10.96] [6.75, 8.18]
Est. and meas. error, EB [2.97, 24.14] [2.24 , 19.31] [1.99, 14.46]
@@
From a policy perspective it is of interest whether the ranking between these links changes. Therefore, Table 2 shows the $95\%$ confidence intervals and credible sets for the difference between link 1 and link 2, and the difference between link 2 and link 3. When we consider estimation error, either separately or combined with measurement error, we fail to reject that these differences are zero and hence lose the ability to rank these three links. When we consider only measurement error, we can reject that these differences are zero. Additional discussion and analyses can be found in Appendices E.2 and E.3.
@@
Only est. error [-6.16, 7.94] [-3.86, 8.32]
Only meas. error, EB [0.39, 5.25] [0.01, 3.31]
Est. and meas. error, EB [-408.36, 411.18] [-396.06, 397.33]
@@ 6 Conclusion
In this paper, I provide an econometric framework for examining the effect of parameter uncertainty and measurement error for an important class of quantitative spatial models. This setting differs from conventional measurement error models, because the object of interest directly depends on the correctly measured data rather than on the correctly measured distribution of the data. I take an empirical Bayes approach to uncertainty quantification and show how to obtain credible sets for the counterfactual change variables of interest. The proposed method accounts for the fact that the structural parameter is often estimated using the noisy data. I discuss the implications of considering measurement error and estimation error for applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For both papers, I find substantial uncertainty in important economic quantities, which highlights the importance of uncertainty quantification.
@@ A.1 Assumptions
The two leading classes of quantitative spatial models I will consider are invertible models and exact hat algebra models. In these models, the following two assumptions on the equilibrium conditions $f$ hold:
@@ A.2 Invertible Models
There is a subset of quantitative spatial models in which there is a one-to-one mapping between the exogenous variables and the endogenous variables. For those models, we are able to invert the model and identify all the exogenous variables from only knowing the endogenous observables. Such models are called invertible in the quantitative spatial literature (Redding and Rossi-Hansberg, 2017). In this case, Assumption 4 will hold with uniqueness. In particular, the exogenous unobservables $X_{U}$ can be recovered from the observables $\left(X_{O},N_{O}\right)$, and there are no endogenous unobservables, so $N_{U}=\left\{\right\}$. One could argue that $X_{U}$ also should be an empty set because we can always recover it, but here I will consider the exogenous unobservables to be variables that are not directly observed in the data.
@@ A.3 Exact Hat Algebra Models
Another important subset of quantitative spatial models contains exact hat algebra models (Costinot and Rodríguez-Clare, 2014). For these models, at least one of $X_{U}$ and $N_{U}$ is non-empty and cannot be recovered from the observables $X_{O}$ and $N_{O}$. We therefore can never identify the counterfactual levels and hence need to shift our focus from levels to changes.
@@ B.1 Counterfactual Change Variable System of Equations for Armington
Define the counterfactual level of a scalar variable by $x^{{}^{\prime}}$, so that $x^{{}^{\prime}}=x\hat{x}$. Note that then for this specific counterfactual question, we have $Q_{i}^{{}^{\prime}}=Q_{i}$ for all $i=1,...,n$. Hence, we have
@@ B.2 Running Example: Exactly Identified Version of Armington
Then, we can identify the endowments or productivities $Q_{i}$ up to a constant:
@@
Appendix C Alternative method: Adversarial Uncertainty Quantification
The empirical Bayes (EB) approach requires a prior on the true data and a distributional assumption on the measurement errors. A contrasting perspective is taken by what I will call the adversarial approach to uncertainty quantification. This approach borrows insights from the literature on “adversarial perturbations” in machine learning, where small changes to images can lead neural networks to misclassify them with high probability (Moosavi-Dezfooli et al., 2017; Balda, Behboodi, and Mathar, 2019).
C.1 Adversarial Uncertainty Quantification
I transfer these ideas to the setting in this paper and define the perturbed counterfactual function
$\tilde{T}_{\boldsymbol{\eta}}$ $=g_{T}\left(\tilde{D}_{O}+\boldsymbol{\eta}\right),$
where $\tilde{D}_{O}=\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ is the observed data vector and $\boldsymbol{\eta}$ is some perturbation vector. The thought experiment of interest is then: how far are we willing to believe that the observed data is from the true data? That is, denoting with $D_{O}=\left(X_{O},N_{O}\right)$ the true data vector, what are constraints on $\boldsymbol{\eta}$ such that we are almost certain that
$D_{O}\in\left[\tilde{D}_{O}-\boldsymbol{\eta},\tilde{D}_{O}+\boldsymbol{\eta}\right].$
Then, given these constraints on the perturbation $\boldsymbol{\eta}$, we are trying to find the largest deviations in the counterfactual outcome of interest that an adversary can cause. The solutions depend on what kind of norm we use to constrain $\boldsymbol{\eta}$. I opt to use the sup-norm, which corresponds to bounding the adversarial perturbations individually. [^6] To bound the adversarial perturbations jointly, one can use the $L_{2}$-norm. Then, using a linear approximation, the perturbations that an adversary would choose would not depend on the signs of the elements of the Jacobian $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)$, but rather on the eigenvector that corresponds to the largest eigenvalue of $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)$. The relevant problems to solve then are:
$\boldsymbol{\eta}_{\min}$ $=\arg\max_{\boldsymbol{\eta}:\left|\boldsymbol{\eta}_{j}\right|\leq\boldsymbol{\epsilon}_{j}\forall j}\tilde{T}_{\boldsymbol{0}}-\tilde{T}_{\boldsymbol{\eta}}$
$\boldsymbol{\eta}_{\max}$ $=\arg\max_{\boldsymbol{\eta}:\left|\boldsymbol{\eta}_{j}\right|\leq\boldsymbol{\epsilon}_{j}\forall j}\tilde{T}_{\boldsymbol{\eta}}-\tilde{T}_{\boldsymbol{0}}.$
If the data is low-dimensional, we can brute-force solve these problems. However, as the dimension of the data increases, this becomes increasingly hard. We can use a linear approximation to obtain the following proposition:
Proposition 2.
If we assume that $T$ is approximately linear in the data, and hence $\tilde{T}_{\boldsymbol{\eta}}$ is approximately linear in $\tilde{D}_{O}$ on $\left[\tilde{D}_{O}-\boldsymbol{\epsilon},\tilde{D}_{O}+\boldsymbol{\epsilon}\right]$, we have
$\boldsymbol{\eta}_{\min}$ $\approx-\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}$
$\boldsymbol{\eta}_{\max}$ $\approx\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}},$
where $\odot$ denotes element-wise multiplication, the function $\mathrm{sign}\left(\cdot\right)$ returns the signs of each element of a vector, and $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)=\frac{\partial g_{T}}{\partial D_{O}}|_{\tilde{D}_{O}}$ denotes the Jacobian of $g_{T}$ at $\tilde{D}_{O}$.
The proof can be found in Appendix G. This first-order adversarial approach is crude, but it is inexpensive to compute and still provides a useful lower bound. Furthermore, we can get a sense of how close it is to the global bound by comparing the predicted values $\tilde{T}\pm\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)\left\{\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}\right\}$ and the actual values $\tilde{T}_{\pm\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}}$ at the first order bounds.
Remark 6.
When we want to simultaneously account for measurement error and estimation error, we have
$\mathbb{P}\left[\hat{k}^{E}\in\left[g_{L_{\hat{k}^{E}}^{\beta}}\left(\tilde{D}_{O}-\boldsymbol{\epsilon}\odot\mathrm{sign}\frac{\partial g_{L_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}\right),g_{U_{\hat{k}^{E}}^{\beta}}\left(\tilde{D}_{O}+\boldsymbol{\epsilon}\odot\mathrm{sign}\frac{\partial g_{U_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}\right)\right]\right]\geq 1-\beta.$
C.2 How to Choose the Bounds on $\boldsymbol{\eta}$
An important question for the adversarial approach to uncertainty quantification is how to choose $\boldsymbol{\epsilon}$. In the applications, I will use the $10\%$ confidence interval of the calibrated measurement error. So if for example we consider normal measurement error with mean 00 and variance $\varsigma^{2}$, the $10\%$ confidence interval will be $\left[-0.13\varsigma,0.13\varsigma\right]$. Researchers could also be agnostic about how large of a confidence interval for the measurement error should be used to choose $\boldsymbol{\epsilon}$. They could then plot the resulting adversarial bounds against the sizes of these confidence intervals and draw a vertical line where they think a valid choice is. I will illustrate such a sensitivity analysis plot for both my applications in Appendices D and E.
Appendix D Details for Application Adao, Costinot, and Donaldson (2017)
D.1 Model Details
@@
$X_{O}$ =$=\left\{Q_{i,t}\right\}$
@@
The counterfactual question of interest is what the change in China’s welfare is due to joining the WTO. This question is modeled by choosing the counterfactual changes in trade costs, $\hat{\tau}_{ij,t}$, such that Chinese trade costs are brought back to their 1995 levels:
@@
Welfare is then defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\hat{\tau}_{ij,t}\tau_{ij,t}\right\}$. These changes in China’s welfare $\left\{\hat{W}_{\mathrm{China},t}^{E}\right\}$ can be obtained from first solving for $\left\{\hat{P}_{i,t}\right\}$ using the system of equations
$\sum_{j}\frac{\exp\left\{\chi_{i}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\tilde{\varepsilon}\log\left(\hat{P}_{i,t}\hat{\tau}_{ij,t}\right)\right\}}{1+\sum_{\ell>1}\exp\left\{\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\tilde{\varepsilon}\log\left(\hat{P}_{\ell,t}\hat{\tau}_{\ell j,t}\right)\right\}}\left\{\hat{P}_{j,t}\hat{Q}_{j,t}Y_{j,t}+\rho_{j,t}\right\}=\hat{P}_{i,t}\hat{Q}_{i,t}Y_{i,t},$
@@
$\hat{W}_{i,t}^{E}=\hat{P}_{i,t}\frac{\sum_{\ell}\left[\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\tilde{\varepsilon}}}{\sum_{\ell}\left[\hat{P}_{\ell,t}\hat{\tau}_{i\ell,t}\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\tilde{\varepsilon}}}.$
D.2 Estimation details
D.2.1 Calibration Procedure
Combining the prior, measurement error model and the fact that we observe two noisy independent observations per time period for each bilateral trade flow, leads to the following hierarchical model:

### v3

@@ 4.3 Results
Having obtained a posterior distribution for the true trade flows given the noisy trade flows, we can now quantify uncertainty about the counterfactual change variables of interest. In Figure 1, I reproduce Figure 3 of Adao, Costinot, and Donaldson (2017), which plots the change in China’s welfare as a result of the China shock for each year in the period 1996-2011, and include $95\%$ intervals. Note that a value of larger than one implies that China’s welfare increased due to the China shock. I plot three different intervals.
The first only considers estimation error and hence assumes the data are perfectly measured. It is constructed using code provided by the authors, which matches the discussion in Section 3.1 and samples from the normal distribution with mean and variance equal to the GMM estimator for the trade elasticity $\varepsilon$ and its sampling variance, respectively. The resulting intervals are small for the period before the year 2000, and then slowly become wider. For the period 2003-2011 it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare. The second region considers only measurement error and no estimation error in $\varepsilon$. The resulting interval is slightly wider than the interval based solely on estimation error, especially in the first few years. Again, for the period 2003-2011 it is unclear whether joining the WTO had a positive or negative effect on China’s welfare.
Finally, the third region combines estimation error and measurement error and follows Algorithm 1. The resulting bounds seem to be reasonable compositions of the bounds considering only estimation error or only measurement error. Now already starting in 2002 it becomes ambiguous whether the China shock had a positive or negative effect.
@@
We can learn more from the posterior distributions than just intervals that cover $\left\{\hat{W}_{\mathrm{China},t}\right\}$. It might for example be of interest what the probability is that the welfare in China increased as a result of the China shock, which is discussed in Appendix C.3. There, I also perform some additional analyses to check the robustness of these results.
@@ 5.1 Model and Counterfactual Question of Interest
The empirical application in Allen and Arkolakis (2022) aims to estimate the returns on investment for all highway segments of the US Interstate Highway network. The authors do so by introducing an economic geography model and calculating what happens to welfare after a $1\%$ improvement to all highway links. Combining these counterfactual welfare changes with how many lane-miles must be added in order to achieve the $1\%$ improvement, they find the highway segments with the greatest return on investment.
This exercise only requires data on incomes and traffic flows of the $n$ locations and knowledge of four structural model parameters. Three of these parameters are taken from the literature and are assumed to have no uncertainty around them. The fourth, which is the congestion elasticity $\delta$, is estimated using the noisily measured traffic flow data. The details of the model can be found in Appendix D.1, but the key relation is the one that maps the log average annual daily traffic (AADT) flows $\left\{\log F_{ij}\right\}$ to the change in welfare $\hat{W}$, which is
$\hat{W}=g_{\hat{W}}\left(\left\{\log F_{ij}\right\};\delta\right)$
for a known function $g_{\hat{W}}:\mathbb{R}_{+}^{n\left(n-1\right)}\rightarrow\mathbb{R}$.
@@ 5.2 Measurement Error Model and Prior
Towards uncertainty quantification, I will add log-normal measurement error to non-zero traffic flows. Musunuru and Porter (2019) estimates that the measurement error variance of the logarithm of the average annual daily traffic (AADT) flows, which is exactly the data that Allen and Arkolakis (2022) uses, is between 0.05 and 0.20. To obtain a lower bound on uncertainty, I will use a uniform measurement error variance of $0.05$. I follow Musunuru and Porter (2019) by adding independent and identically distributed measurement error shocks to log traffic flows. Hence, this fits the setting from Section 3.4 and the prior distribution and measurement error model are as in Equation (9).
Using $\hat{\varsigma}^{2}=0.05$, I estimate a prior variance of $\hat{s}^{2}=0.101$. This results in the following estimated posterior distribution for the true traffic flow between country $i$ and $j$, $\log F_{ij}$, given its noisy version $\log\tilde{F}_{ij}$, for $i,j=1,...,n$:
$\mathcal{N}\left(0.669\cdot\log\tilde{F}_{ij}+0.331\cdot\left\{\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\},0.033\right).$
@@ 5.3 Results
The counterfactual question of interest is which links have the highest return on investment, and the authors of Allen and Arkolakis (2022) report the top ten links. For exposition, I will focus my analysis on the three best performing links. Similarly to the setting in Section 4, I consider scenarios with only estimation error, only measurement error, and both estimation and measurement error.
Concerning estimation error, I follow the discussion in Section 3.1 and sample from the normal distribution with mean and variance equal to the IV estimator for the congestion elasticity $\delta$ and its squared clustered standard error, respectively. Table 1 shows the $95\%$ equal-tailed intervals for the top three links. The intervals that consider just measurement error or estimation error are of similar order of magnitude, and the intervals that combines them seem a sensible composition.
@@
Only est. error [8.33, 11.47] [7.02, 10.76] [5.05, 8.57]
Only meas. error, EB [8.69, 14.15] [7.31, 10.83] [6.78, 8.18]
Est. and meas. error, EB [7.86, 14.89] [6.60, 11.32] [5.30, 8.90]
@@
From a policy perspective it is of interest whether the ranking between these links can change due to estimation and measurement error. Therefore, Table 2 shows the $95\%$ intervals for the difference between link 1 and link 2, and the difference between link 2 and link 3. It follows that the rankings are generally robust against estimation error and measurement error. Additional discussion and analyses can be found in Appendices D.2 and D.3.
@@
Only est. error [0.61, 1.29] [1.96, 2.25]
Only meas. error, EB [0.38, 5.39] [-0.05, 3.27]
Est. and meas. error, EB [0.38, 5.66] [0.02, 3.49]
@@ 6 Conclusion
In this paper, I provide an econometric framework for examining the effect of parameter uncertainty and measurement error for an important class of quantitative trade and spatial models. This setting differs from conventional measurement error models, because the object of interest directly depends on the correctly measured data rather than on the correctly measured distribution of the data. I take an empirical Bayes approach to uncertainty quantification and show how to quantify uncertainty about the counterfactual change variables of interest. The proposed method accounts for the fact that the structural parameter is often estimated using the noisy data. I discuss the implications of considering measurement error and estimation error for settings in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For both papers, I find substantial uncertainty in important economic quantities, which highlights the importance of uncertainty quantification.
@@ A.1 Assumptions
The two leading classes of quantitative trade and spatial models I will consider are invertible models and exact hat algebra models. In these models, the following two assumptions on the equilibrium conditions $f$ hold:
@@ A.2 Invertible Models
There is a subset of quantitative trade and spatial models in which there is a one-to-one mapping between the exogenous variables and the endogenous variables. For those models, we are able to invert the model and identify all the exogenous variables from only knowing the endogenous observables. Such models are called invertible in the quantitative spatial literature (Redding and Rossi-Hansberg, 2017). In this case, Assumption 4 will hold with uniqueness. In particular, the exogenous unobservables $X_{U}$ can be recovered from the observables $\left(X_{O},N_{O}\right)$, and there are no endogenous unobservables, so $N_{U}=\left\{\right\}$. One could argue that $X_{U}$ also should be an empty set because we can always recover it, but here I will consider the exogenous unobservables to be variables that are not directly observed in the data.
@@ A.3 Exact Hat Algebra Models
Another important subset of quantitative trade and spatial models contains exact hat algebra models (Dekle, Eaton, and Kortum, 2008; Costinot and Rodríguez-Clare, 2014). For these models, at least one of $X_{U}$ and $N_{U}$ is non-empty and cannot be recovered from the observables $X_{O}$ and $N_{O}$. We therefore can never identify the counterfactual levels and hence need to shift our focus from levels to changes.
@@ B.1 Counterfactual Change Variable System of Equations for Armington
Define the counterfactual level of a scalar variable by $x^{{}^{\prime}}$, so that $x^{{}^{\prime}}=x\hat{x}$. Note that then for the counterfactual question of perturbing the trade costs, we have $Q_{i}^{{}^{\prime}}=Q_{i}$ for all $i=1,...,n$. Hence, we have
@@ B.2 Running Example: Exactly Identified Version of Armington
Then, we can identify the endowments $Q_{i}$ up to a constant:
@@
Appendix C Details for Application Adao, Costinot, and Donaldson (2017)
C.1 Model Details
@@
$X_{O}$ =($=\left(\left\{Q_{i,t}\right\},\left\{\rho_{i,t}\right\}\right)$
@@
The counterfactual question of interest is what the change in China’s welfare is due to joining the WTO. This question is modeled by choosing the counterfactual changes in trade costs, $\left\{\hat{\tau}_{ij,t}\right\}$, such that Chinese trade costs are brought back to their 1995 levels:
@@
Welfare is then defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\hat{\tau}_{ij,t}\tau_{ij,t}\right\}$. These changes in China’s welfare $\left\{\hat{W}_{\mathrm{China},t}\right\}$ can be obtained from first solving for $\left\{\hat{P}_{i,t}\right\}$ using the system of equations
$\sum_{j}\frac{\exp\left\{\chi_{i}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\varepsilon\log\left(\hat{P}_{i,t}\hat{\tau}_{ij,t}\right)\right\}}{1+\sum_{\ell>1}\exp\left\{\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\varepsilon\log\left(\hat{P}_{\ell,t}\hat{\tau}_{\ell j,t}\right)\right\}}\left\{\hat{P}_{j,t}Y_{j,t}+\rho_{j,t}\right\}=\hat{P}_{i,t}Y_{i,t},$
@@
$\hat{W}_{i,t}=\hat{P}_{i,t}\frac{\sum_{\ell}\left[\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\varepsilon}}{\sum_{\ell}\left[\hat{P}_{\ell,t}\hat{\tau}_{i\ell,t}\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\varepsilon}}-1.$
C.2 Estimation details
C.2.1 Calibration Procedure
Combining the prior distribution, measurement error model and the fact that we observe two noisy independent observations per time period for each bilateral trade flow, leads to the following hierarchical model: